OpenStudy (anonymous):

What is the value of the expression 1/5 To the power of -5

OpenStudy (jdoe0001):

\(\bf \left(\cfrac{1}{5}\right)^{-5}\quad ?\)

OpenStudy (anonymous):



OpenStudy (jdoe0001):

keep in mind that \(\bf a^{-{\color{red} n}} = \cfrac{1}{a^{\color{red} n}}\) so \(\bf \cfrac{1}{5^{-5}}\implies \cfrac{1}{\frac{1}{5^5}}\implies \large \cfrac{\frac{1}{1}}{\frac{1}{5^5}}\\ \quad \\ \textit{keep in mind that }\cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\cdot \cfrac{d}{c}\) so what would that give you?

OpenStudy (anonymous):

25 ?

OpenStudy (jdoe0001):

25?

OpenStudy (jdoe0001):

well... how did you get 25 anyway?

OpenStudy (anonymous):

let me see.... Ok 3125

OpenStudy (jdoe0001):

3125 \(\checkmark\)

OpenStudy (anonymous):

Thank you.